# Offline changepoint localization using a matrix of conformal p-values

### v4

@@ 4.1 Hypothesis testing: generating left and right p-values
We now discuss how we can consolidate the matrix of p-values from the MCP algorithm into two p-values for each $\mathcal{H}_{0t}$, which we call the left and right p-values. This consolidation can be done using an empirical test. We can use simulations to construct confidence sets for the changepoint by estimating the level $1-\alpha$ quantile of $W_{t}^{(0)}$ and $W_{t}^{(1)}$ under the null hypothesis that there is no changepoint. We use Algorithm 2 to construct left and right p-values. We also detail an alternative method based on an asymptotic KS test or permutation test in Appendix C, which can be used to speed up our algorithm when $t$ and ntn-t are large.
Input: $(W_{t}^{(0)},W_{t}^{(1)})$ (discrepancy scores), $B\in\mathbb{N}$ (sample size)
@@
1
2for $b\in[B]$ do
3 $(X_{t}^{(b)})_{t=1}^{n}\sim\mathrm{Unif}([0,1]^{n})$
4  Compute $(W_{t}^{(0,b)})_{t=1}^{n-1}$ and $(W_{t}^{(1,b)})_{t=1}^{n-1}$ using algorithm 1 to algorithm 1 of Algorithm 1 on simulated data $(X_{t}^{(b)})_{t=1}^{n}$
5 end for
6
7Draw $\theta_{0},\,\theta_{1}\sim\mathrm{Unif}(0,1)$
[…]
@@ 4.2 Combining p-values and constructing confidence sets
Notice that the left and right p-values are independent if the score functions $s_{rt}^{(0)}$ and $s_{rt}^{(1)}$ have no dependence on its last argument; we call such score functions non-adaptive. Intuitively, a non-adaptive left score function $s_{t}^{(0)}$ is one which does not use data to the right of $t$. On the other hand, it’s clear that the left and right p-values won’t necessarily be independent if the score functions are adaptive ($s_{t}^{(0)}$ or $s_{t}^{(1)}$ depend nontrivially on their last argument).
  • Minimum (requires independence): The choice $p_{t}=1-(1-\min\{p_{t}^{\mathrm{left}},p_{t}^{\mathrm{right}}\})^{2}$ is a powerful combining method for change detection, since away from the changepoint, we would expect either the left or right p-value to be small. Under the null hypothesis $\mathcal{H}_{0t}$, we would expect $p_{t}^{\mathrm{left}}$ and $p_{t}^{\mathrm{right}}$ to be independent uniform random variables, so that $p_{t}$ is distribution-free and uniformly distributed. Note that one can use Fisher’s method ([8]) given by $p_{t}=F_{\chi^{2}_{4}}(-2\log(p_{t}^{\mathrm{left}})-2\log(p_{t}^{\mathrm{right}}))$; however, we have observed the minimum method to work better in practice.
@@ 5 How should the score function be chosen?
$T_{n}[s]=\frac{1}{n}\sum_{i=1}^{n}(\mathbf{1}_{s(X_{i})>s(X_{n+1})}+\theta_{i}\,\mathbf{1}_{s(X_{i})=s(X_{n+1})}),$
@@
Note that the proof of Theorem 5.1 (given in Appendix E) also shows a stronger result; define the conditional power
@@
In this sense, the likelihood ratio score yields the uniformly most powerful conformal test. A similar result can be shown without randomizing using the $\theta_{i}$, assuming that the scores are almost surely distinct. Appendix F presents an analogous result for e-values.
@@
We give several additional methods for choosing the score function in Appendix H. Finally, we demonstrate the performance of our algorithms through simulations.
@@ 6 Experiments
https://github.com/sanjitdp/conformal-change-localization.
@@
All experiments are run using the CPU on our personal MacBook Pro with an M1 Pro processor; our algorithm takes under 10 seconds to run in each experiment. For further experiments demonstrating the performance of the MCP algorithm in localizing a Gaussian mean change and visualizing the matrix of conformal p-values, see Appendix A.
6.1 MNIST digit change
We first consider a simulation of a digit change from the MNIST handwritten digit dataset [6]. Suppose that we observe data of length $n=1000$ and the changepoint is $\xi=400$; the pre-change class $\mathbb{P}_{0}$ consists of i.i.d. draws from the set of handwritten “3” digits and the post-change class $\mathbb{P}_{1}$ consists of i.i.d. draws from the set of handwritten “7” digits (Figure 3).
As described in Section 5, we can use a pretrained multi-class classifier to estimate the likelihood ratio using the data before and after $t$ (for each $t\in[n-1]$). In this case, we use a simple convolutional neural network; see Appendix G for details about how the network was trained.

### v5

@@ 4.1 Hypothesis testing: generating left and right p-values
We now discuss how we can consolidate the matrix of p-values from the MCP algorithm into two p-values for each $\mathcal{H}_{0t}$, which we call the left and right p-values. This consolidation can be done using an empirical test. We can use simulations to construct confidence sets for the changepoint by estimating the level $1-\alpha$ quantile of $W_{t}^{(0)}$ and $W_{t}^{(1)}$ under the null hypothesis that there is no changepoint. We use Algorithm 2 to construct left and right p-values. We also detail an alternative method based on an asymptotic KS test or permutation test in Appendix D, which can be used to speed up our algorithm when $t$ and ntn-t are large. Performing the empirical test technically adds $O(Bn)$ to the time complexity, but is typically very fast in practice due to existence of pre-computed quantiles for the KS test in software packages.
Input: $(W_{t}^{(0)},W_{t}^{(1)})$ (discrepancy scores), $B\in\mathbb{N}$ (user-chosen sample size)
@@
for $b\in[B]$ do
 $(X_{t}^{(b)})_{t=1}^{n}\sim\mathrm{Unif}([0,1]^{n})$
  Compute $(W_{t}^{(0,b)})_{t=1}^{n}$ and $(W_{t}^{(1,b)})_{t=1}^{n}$ using algorithm 1 to algorithm 1 of Algorithm 1 on simulated data $(X_{t}^{(b)})_{t=1}^{n}$
end for
Draw $\theta_{0},\,\theta_{1}\sim\mathrm{Unif}(0,1)$
$p_{t}^{\mathrm{left}}\leftarrow\frac{1}{B+1}\left(\theta_{0}+\sum_{b=1}^{B}(\mathbf{1}_{W_{t}^{(0,b)}>W_{t}^{(0)}}+\theta_{0}\,\mathbf{1}_{W_{t}^{(0,b)}=W_{t}^{(0)}})\right)$
$p_{t}^{\mathrm{right}}\leftarrow\frac{1}{B+1}\left(\theta_{1}+\sum_{b=1}^{B}(\mathbf{1}_{W_{t}^{(1,b)}>W_{t}^{(1)}}+\theta_{1}\,\mathbf{1}_{W_{t}^{(1,b)}=W_{t}^{(1)}})\right)$
@@ 4.2 Combining p-values and constructing confidence sets
Notice that the left and right p-values are independent if the score functions $s_{rt}^{(0)}$ and $s_{rt}^{(1)}$ have no dependence on its last argument; we call such score functions non-adaptive. Intuitively, a non-adaptive left score function $s_{rt}^{(0)}$ is one which does not use data to the right of $t$. On the other hand, it’s clear that the left and right p-values won’t necessarily be independent if the score functions are adaptive ($s_{rt}^{(0)}$ or $s_{rt}^{(1)}$ depend nontrivially on their last argument).
  • Minimum (requires independence): The choice $p_{t}=1-(1-\min\{p_{t}^{\mathrm{left}},\,p_{t}^{\mathrm{right}}\})^{2}$ is a powerful combining method for change detection, since away from the changepoint, we would expect either the left or right p-value to be small. Under the null hypothesis $\mathcal{H}_{0t}$, we would expect $p_{t}^{\mathrm{left}}$ and $p_{t}^{\mathrm{right}}$ to be independent uniform random variables, so that $p_{t}$ is distribution-free and uniformly distributed. Note that one can use Fisher’s method ([fisher1928statistical]) given by $p_{t}=F_{\chi^{2}_{4}}(-2\log(p_{t}^{\mathrm{left}})-2\log(p_{t}^{\mathrm{right}}))$; however, we have observed the minimum method to work better in practice.
@@
4.3 Theoretical results
We begin with the following finite-sample coverage guarantee for the MCP algorithm (Algorithm 1).
Theorem 4.1 (Coverage guarantee (empirical test)).
$\mathbb{P}\left(\xi\in\mathcal{C}_{1-\alpha}\right)\geq 1-\alpha.$
In particular, note that if there is no changepoint, then $\xi=n$, and the confidence set will contain $n$ with probability at least $1-\alpha$. We will see in experiments that when $\xi\neq n$, the confidence sets typically omit $n$.
Recently, [bhattacharya2025theoretical] showed that in the parametric setting and under relatively general triangular array assumptions, the relative length of the confidence set produced by the MCP algorithm converges to zero as $n\to\infty$. Under the same assumptions, that paper also showed consistency of the point estimator $\hat{\xi}\coloneqq\operatorname*{\arg\max}_{t\in[n]}\;p_{t}$ to the true changepoint $\xi$ at the rate $o_{p}(n^{1/4})$.
@@ 5 How should the score function be chosen?
$T_{n}[s]=\frac{1}{n}\sum_{i=1}^{n}(\mathbf{1}_{s(X_{i})<s(X_{n+1})}+\theta_{i}\,\mathbf{1}_{s(X_{i})=s(X_{n+1})}),$
@@
Define the test $\phi_{\alpha}[s]\coloneqq\mathbf{1}_{T_{n}[s]\geq t_{n,\,1-\alpha}}$ for some fixed threshold $t_{n,\,1-\alpha}\geq 0$ and let $\alpha$ denote the level of $\phi_{\alpha}[s^{*}]$ ($\alpha=\mathbb{E}[\phi_{\alpha}[s^{*}]]$ when $Q=R$). Then for any level $\alpha$ test $\phi$ of $\mathcal{H}_{0}$ against $\mathcal{H}_{1}$, it holds that
$\mathbb{E}[\phi_{\alpha}[s^{*}]]\geq\mathbb{E}[\phi].$
In particular, it follows that
$\mathbb{E}[\phi_{\alpha}[s^{*}]]\geq\mathbb{E}[\phi_{\alpha}[s]]$
for any choice of $s$.
Note that the proof of Theorem 5.1 (given in Appendix F) also shows a stronger result; define the conditional power
@@
In this sense, the likelihood ratio score yields the uniformly most powerful conformal test. A similar result can be shown without randomizing using the $\theta_{i}$, assuming that the scores are almost surely distinct. Appendix G presents an analogous result for e-values.
@@
By precomputing the most common class for the prefix and postfix arrays, this score function does not incur any additional computational cost over using the oracle likelihood ratio score function. We give several additional methods for choosing the score function in Appendix I. Finally, we demonstrate the performance of our algorithms through simulations.
6 Extension to multiple changepoints
In this section, we describe how the MCP algorithm (Algorithm 1) can be used to localize multiple changepoints. Suppose we are interested in a model where $(X_{\xi_{k}+1},\dots,X_{\xi_{k+1}})\sim\mathbb{P}_{k}$ for $0\leq k\leq K$, where $0=\xi_{0}<\xi_{1}<\dots<\xi_{K+1}=n$. Since we would like to apply the MCP algorithm, we assume that $\mathbb{P}_{k}$ is exchangeable for all $0\leq k\leq K$ (the data distribution between changepoints is exchangeable). Now, we would like to recover a confidence set which covers all of the changepoints $\{\xi_{k}\}_{k=1}^{K}$ with high probability.
To achieve this goal, we consider the popular kernel changepoint detection (KCPD) framework for finding point estimators of multiple changepoints [harchaoui2007retrospective, arlot2019kernel]. Under mild detectability and minimum separation assumptions, denoting the minimum separation between by $\ell$ and fixing any $\epsilon>0$, [diaz2025consistent] recently showed that the KCPD estimator is consistent in the sense that the number of estimated changepoints is $K$ and the estimated changepoints are all ($\epsilon\ell$)-close to the true changepoints, with probability tending to one.
Now, suppose we are on the high-probability event where the estimated number of changepoints is $K$ and the estimated changepoints are all ($\ell/2$)-close to the true changepoints; denote the estimated changepoints by $(\hat{\xi}_{k})_{k=1}^{K}$. Let $t_{k}\coloneqq(\hat{\xi}_{k}+\hat{\xi}_{k-1})/2$ for $k\in[K+1]$, defining $\hat{\xi}_{0}\coloneqq 0$ and $\hat{\xi}_{K+1}\coloneqq n$. Then, the segments of data $(X_{t_{k}},\,X_{t_{k+1}})$ have exactly one changepoint contained within them for $k\in[K]$, such that the data to the left and right of the true changepoint is exchangeable. Thus, the MCP algorithm applied separately to each of the intervals $(X_{t_{k}},\,X_{t_{k+1}})$ produces a $1-\alpha$ confidence set $\mathcal{C}^{(k)}_{1-\alpha}$ for $\xi_{k}$.
Putting the above considerations together, we can produce a heuristic asymptotic $1-\alpha$ confidence set for $\xi_{k}$ using the MCP algorithm as a wrapper over the KCPD algorithm. We exemplify this method to localize multiple Gaussian mean changes. In this simulation, we choose $n=1500$ with $K=4$ and $(\xi_{1},\,\xi_{2},\,\xi_{3},\,\xi_{4})=(150,\,500,\,820,\,1100)$. Here, we choose $\mathbb{P}_{k}=\mathcal{N}(\mu_{k},\,1)$ with $(\mu_{0},\,\mu_{1},\,\mu_{2},\,\mu_{3},\,\mu_{4})=(-1,\,1/2,\,3/2,\,-2,\,-1)$. We depict the input data in Figure 3.
Figure 3: Sample of multiple Gaussian mean change data.
We apply the KCPD algorithm as described in [diaz2025consistent] using the squared exponential kernel
$k(z,z^{\prime})=\exp\left(\frac{(z-z^{\prime})^{2}}{2\sigma^{2}}\right),$
where $\sigma>0$ is chosen to be the median of $(\lvert X_{i}-X_{j}\rvert)_{1\leq i<j\leq n}$. Then, we apply the MCP algorithm on the intervals $(X_{t_{k}},\,X_{t_{k+1}})$ for $k\in[K]$ using the oracle likelihood ratio score function as described in Section 5; we depict the results in Figure 4.
The estimated changepoints from KCPD are (150, 497, 820, 1091)(150,\,497,\,820,\,1091). Now the MCP confidence set for $\xi_{1}$ is $[111,163]$, the confidence set for $\xi_{2}$ is $[444,541]$, the confidence set for $\xi_{3}$ is $[807,846]$, and the confidence set for $\xi_{4}$ is $[1072,1163]$. In particular, once the KCPD algorithm isolates each changepoint, the MCP algorithm is able to produce confidence sets for each changepoint separately. This is an instance of the popular “Isolate-Detect” paradigm in changepoint localization [anastasiou2022detecting], which consists of isolating each changepoint to lie in an interval and then applying a single-changepoint algorithm.
Figure 4: p-values for multiple Gaussian mean changes using MCP as a wrapper on KCPD. The dashed red lines indicate the true changepoints and the dashed yellow lines indicate the point estimators from KCPD. The regions on the horizontal axis where the p-values lie above the dotted green line ($\alpha=0.05$) correspond to our heuristic 95% confidence sets for each changepoint.
@@ 7 Experiments
All experiments are run using the CPU on our personal MacBook Pro with an M1 Pro processor; our algorithm takes under 10 seconds to run in each experiment.
7.1 Gaussian mean change
We begin by discussing our test for $\mathcal{H}_{0n}$ in Algorithm 1. When we test $\mathcal{H}_{0n}$ at level $\alpha_{0}=0.01$ using the identity score function and only the forward p-values, we estimate over 1000 simulations (with no changepoint) that the Type I error rate of our test for $\mathcal{H}_{0n}$ is $0.009$, which aligns closely with the theoretical Type I error of $\alpha_{0}=0.01$. In these cases (when there is no changepoint), the resulting confidence set is sometimes much wider, but the above discussion shows that it usually correctly contains $n$; we expand on these findings in Appendix E. On the other hand, over 1000 simulations (with a changepoint), the test for $\mathcal{H}_{0n}$ was always able to detect a deviation from exchangeability and had an empirical power of 1.000.
In all future experiments, the dataset contains a changepoint and the test for $\mathcal{H}_{0n}$ correctly rejects the null hypothesis; therefore, $n$ is not included in the resulting confidence sets. Suppose that we do not have access to the oracle score function. Then, as described in Section 5, we can use a kernel density estimator to estimate the likelihood ratio using the data before and after $t$ (for each $t\in[n-1]$), using a Gaussian kernel. We choose bandwidth $10^{-1}$ when $t\in\{1,n-1\}$ and bandwidth $r^{-1/5}$ otherwise, where $r$ is the number of samples used to learn each KDE; this is known as Scott’s rule (due to [scott1979optimal]). Then, we repeat the process described above to compute the left and right p-values. However, due to dependence of the left and right p-values, we use the Bonferroni combining rule, plotting the resulting p-values $(p_{t})_{t=1}^{n-1}$ in Figure 5b.
(a) oracle score function family
(b) KDE learned score function family
Figure 5: p-values for a Gaussian mean change at $\xi=400$ (a single run). The dashed red line indicates the true changepoint, and the region on the horizontal axis where the p-values lie above the dotted green line ($\alpha=0.05$) corresponds to our 95% confidence set. The point estimators are (a) $\hat{\xi}=409$ and (b) $\hat{\xi}=380$, both of which are close to the true $\xi=400$.
Next, in Figure 6, we plot the resulting p-values when the true changepoint is closer to the edge of the dataset. We observe that the size of the confidence set and quality of the point estimator do not suffer too much even when the changepoint is very close to the edge of the dataset. By symmetry, we expect similar results to hold when the true changepoint is in the right half of the data.
(a) oracle score function family
(b) KDE learned score function family
Figure 6: p-values for a Gaussian mean change at (a) $\xi=50$ and (b) $\xi=200$ (a single run). The dashed red line indicates the true changepoint, and the region on the horizontal axis where the p-values lie above the dotted green line ($\alpha=0.05$) corresponds to our 95% confidence set. The point estimators are (a) $\hat{\xi}=14$ and (b) $\hat{\xi}=189$, both of which are relatively close to the true changepoints. The sizes of the confidence sets are (a) 60 and (b) 53.
7.1.1 Frequentist operating characteristics of MCP
Running the same simulation 200 times for a change from $\mathcal{N}(-1,1)$ to $\mathcal{N}(1,1)$ at $\xi=400$ with $n=1000$ samples, we plot the empirical power of our test for $\mathcal{H}_{0t}$ against various values of $t\neq\xi=400$ in Figure 7. As expected, we observe that the power of the test increases as we move further away from the true changepoint.
(a) oracle score function family
(b) KDE learned score function family
Figure 7: Empirical power of the MCP test for $\mathcal{H}_{0t}$ against various values of $t\neq\xi=400$, using (a) the oracle likelihood ratio score function (averaged over 200 runs) and (b) the KDE score function (averaged over 10 runs).
Additionally, we report the empirical average width and coverage of our confidence sets at various levels, as well as the bias and mean absolute deviation of our point estimator $\hat{\xi}=\arg\max_{t\in[n]}\;p_{t}$, in Table 1. In Table 1, we also compare the performance of the MCP algorithm on datasets with varying coverage, signal strength and sample size.
$n$ $\delta$ Target $\alpha$ Avg. width Empirical coverage Bias Mean abs. dev.
100 2.0 0.05 $22.29_{\mathrm{se:}\,6.9}$ $0.93_{\pm 0.03}$ $0.14_{\mathrm{se:}\,4.7}$ $3.50_{\mathrm{se:}\,3.1}$
0.5 $6.63_{\mathrm{se:}\,4.9}$ $0.50_{\pm 0.07}$
1.0 0.05 $31.41_{\mathrm{se:}\,10.1}$ $0.47_{\pm 0.03}$ $0.33_{\mathrm{se:}\,6.6}$ $5.10_{\mathrm{se:}\,4.6}$
0.5 $9.03_{\mathrm{se:}\,6.4}$ $0.94_{\pm 0.07}$
200 2.0 0.05 $29.18_{\mathrm{se:}\,8.0}$ $0.96_{\pm 0.03}$ $0.50_{\mathrm{se:}\,5.6}$ $4.43_{\mathrm{se:}\,3.6}$
0.5 $8.88_{\mathrm{se:}\,6.6}$ $0.54_{\pm 0.07}$
1.0 0.05 $38.96_{\mathrm{se:}\,11.7}$ $0.97_{\pm 0.03}$ $0.12_{\mathrm{se:}\,7.9}$ $5.96_{\mathrm{se:}\,5.4}$
0.5 $12.12_{\mathrm{se:}\,8.7}$ $0.53_{\pm 0.07}$
500 2.0 0.05 $41.05_{\mathrm{se:}\,12.0}$ $0.93_{\pm 0.03}$ $2.22_{\mathrm{se:}\,8.7}$ $7.02_{\mathrm{se:}\,5.3}$
0.5 $13.04_{\mathrm{se:}\,10.3}$ $0.53_{\pm 0.07}$
1.0 0.05 $55.13_{\mathrm{se:}\,16.7}$ $0.94_{\pm 0.03}$ $0.78_{\mathrm{se:}\,12.6}$ $8.96_{\mathrm{se:}\,8.5}$
0.5 $17.42_{\mathrm{se:}\,13.3}$ $0.50_{\pm 0.07}$
Table 1: Average width and coverage of confidence sets for Gaussian mean change, as well as the bias ($\mathbb{E}[\hat{\xi}-\xi]$) and mean absolute deviation ($\mathbb{E}[\lvert\hat{\xi}-\xi\rvert$]) of the point estimator over 1000 trials. The length of the data is $n$ and we observe a change from $\mathcal{N}(-\delta,1)$ to $\mathcal{N}(\delta,1)$ at $\xi=2n/5$. We run the MCP algorithm using the oracle likelihood ratio score function to construct a $1-\alpha$ confidence set for the changepoint. Note that the empirical coverage is very close to $1-\alpha$, and the bias is very small relative to $n$ (though systematically negative, because the changepoint is in the left half). Binomial errors are reported for empirical coverage (95%) and standard errors are reported for all other values.
We know from Theorem 4.1 that the true coverage probability is at least 50% and 95% respectively for a 50% or 95% confidence interval constructed using our method; this is what we observe empirically as well. Furthermore, we observe that the point estimator is empirically close to the true changepoint; this is expected due to a recent result of [bhattacharya2025theoretical] showing that the point estimator $\hat{\xi}$ is consistent at the rate $o_{p}(n^{1/4})$ in a general triangular array setup. Additionally, we observe in practice that because the changepoint is in the left half of the data, the point estimator tends to be biased towards the left of the true changepoint.
7.1.2 Comparison against baselines
Next, we compare the performance of the MCP algorithm against several baselines for changepoint localization. As discussed in Section 2, there are relatively few existing methods for constructing confidence sets for changepoints, especially in the nonparametric setting. We consider one baseline algorithm from [siegmund1988confidence], which constructs confidence intervals for the changepoint under a (parametric) exponential family assumption by deriving an approximation to the asymptotic distribution of the MLE. We also consider the bootstrap-based algorithm from [cho2022bootstrap], which constructs confidence intervals for a nonparametric univariate mean change. We compare the average width and empirical coverage of our confidence sets against these baselines in Table 2. Here, we run all methods on data of length $n=1000$ with a change from $\mathcal{N}(-1,1)$ to $\mathcal{N}(1,1)$ at $\xi=400$.
Method Avg. width Empirical coverage Bias Mean abs. dev.
MCP (ours) $41.69_{\mathrm{se:}\,0.6}$ $0.96_{\pm 0.03}$ $0.32_{\mathrm{se:}\,9.0}$ $7.02_{\mathrm{se:}\,5.6}$
[siegmund1988confidence] $1.31_{\mathrm{se:}\,0.5}$ $1.00_{\pm 0.03}$ $0.04_{\mathrm{se:}\,0.3}$ $0.15_{\mathrm{se:}\,0.2}$
[cho2022bootstrap] $3.00_{\mathrm{se:}\,0.1}$ $0.94_{\pm 0.03}$ $1.04_{\mathrm{se:}\,0.3}$ $1.05_{\mathrm{se:}\,0.3}$
Table 2: Average width and empirical coverage of confidence sets for Gaussian mean change, as well as the bias ($\mathbb{E}[\hat{\xi}-\xi]$) and mean absolute deviation ($\mathbb{E}[\lvert\hat{\xi}-\xi\rvert$]) of the point estimator over 1000 trials. The length of the data is $n=1000$ and we observe a change from $\mathcal{N}(-1,1)$ to $\mathcal{N}(1,1)$ at $\xi=400$. We run the MCP algorithm using the oracle likelihood ratio score function to construct a 95% confidence set for the changepoint, and compare against the methods of [siegmund1988confidence] and [cho2022bootstrap]. Binomial errors are reported for empirical coverage (95%) and standard errors are reported for all other values.
We observe from Table 2 that the methods of [siegmund1988confidence] and [cho2022bootstrap] do produce sharper confidence sets under their respective modeling assumptions, but our MCP algorithm operates under far more general assumptions. For instance, our method does not require the existence of any moments of the data distribution, and is able to operate on high-dimensional data. To illustrate this point, we consider a change from a Cauchy distribution with location 1-1 to a Cauchy distribution with location 1 at $\xi=400$. The method of [cho2022bootstrap] is also not applicable here, since the Cauchy distribution does not have a finite mean, and [siegmund1988confidence] is misspecified here, since the data is not Gaussian. We show in Table 3 that these methods fare poorly, but our MCP algorithm is able to produce valid confidence sets even in this challenging scenario, even with misspecified score function (i.e. even if we incorrectly assumed the data was Gaussian to design the score, our coverage is maintained).
Method Avg. width Empirical coverage Bias Mean abs. dev.
MCP – oracle score (ours) $53.27_{\mathrm{se:}\,17.4}$ $0.97_{\pm 0.03}$ $0.25_{\mathrm{se:}\,19.8}$ $8.78_{\mathrm{se:}\,10.9}$
MCP – Gaussian score (ours) $70.69_{\mathrm{se:}\,20.1}$ $0.94_{\pm 0.03}$ $3.685_{\mathrm{se:}\,32.2}$ $17.40_{\mathrm{se:}\,20.3}$
[siegmund1988confidence] $1.01_{\mathrm{se:}\,0.6}$ $0.12_{\pm 0.03}$ $31.79_{\mathrm{se:}\,110.9}$ $77.94_{\mathrm{se:}\,84.5}$
[cho2022bootstrap] $181.46_{\mathrm{se:}\,48.1}$ $0.82_{\pm 0.03}$ $23.96_{\mathrm{se:}\,67.7}$ $50.00_{\mathrm{se:}\,53.2}$
Table 3: Average width and empirical coverage of confidence sets for a Cauchy location change over 1000 trials. The length of the data is $n=1000$ and we observe a change from $\mathrm{Cauchy}(-1,1)$ to $\mathrm{Cauchy}(1,1)$ at $\xi=400$. We run the MCP algorithm using both the oracle likelihood ratio score function as well as the misspecified likelihood ratio score function assuming that there is a change from $\mathcal{N}(-1,1)$ to $\mathcal{N}(1,1)$, to construct a 95% confidence set for the changepoint. We compare against the method of [siegmund1988confidence], which produces narrow confidence sets with extremely low coverage, as well as the method of [cho2022bootstrap], which produces wide confidence sets with low coverage in this setting. Binomial errors are reported for empirical coverage (95%) and standard errors are reported for all other values.
Furthermore, as we demonstrate in the following sections, our MCP algorithm is able to operate in many practical settings (e.g., text data, accelerometer data, and image data) where existing methods such as those of [siegmund1988confidence] and [cho2022bootstrap] are not applicable.
7.1.3 Visualizing the matrix of conformal p-values
To give more intuition for the MCP algorithm, we can also visualize the matrix of conformal p-values (MCP) as described in Section 3, in Figure 8. We consider the setting where $n=1000$ and there is a change from $\mathcal{N}(-1,1)$ to $\mathcal{N}(1,1)$ at $\xi=400$. Observe that only when $t=\xi=400$ are the p-values in row $t$ truly uniformly distributed and independent in the left and right halves respectively, as they should be under $\mathcal{H}_{0t}$. For instance, if $t<\xi$, then the right p-values will be invalid and the Kolmogorov-Smirnov statistic $W_{t}^{(1)}$ used in the MCP algorithm (Algorithm 1) will detect their deviation from uniformity.
(a) matrix of conformal p-values (MCP)
(b) validity of conformal p-values